# Deeper or Wider: A Perspective from Optimal Generalization Error with Sobolev Loss

### v3

@@ 4.1 Generalization Error in deep Ritz Method
We initially focus on solving the Poisson equation, as represented by Eq. (6), using the deep Ritz method. This approach utilizes the energy as the loss function during training to guide the neural networks in approximating the solutions. In the context of the Poisson equation, the energy functional encompasses the first derivative of the neural networks. Consequently, the corresponding loss function in deep Ritz methods can be expressed as:
$\mathcal{E}_{D,r}(\phi):=$ $\int_{\widehat{\Omega}}g(\bm{x};\bm{\theta})\,\mathrm{d}\bm{x}+\lambda\int_{\partial\widehat{\Omega}}\phi(\bm{x};\bm{\theta})^{2}\,\mathrm{d}s,$
@@
where $\bm{\theta}$ represents all the parameters in the neural network, and $\lambda$ is a constant to balance the interior and boundary terms. To discretize the loss function, we randomly choose $M_{1}$ points in the interior and $M_{2}$ points on the boundary:
$\mathcal{E}_{S,r}(\phi):=$ $\frac{1}{2M_{1}}\sum_{i=1}^{M_{1}}g(\bm{x}_{i};\bm{\theta})+\frac{\lambda}{M_{2}}\sum_{i=1}^{M_{2}}\phi^{2}(\bm{y}_{i};\bm{\theta}).$
@@
Suppose that $\phi\in\mathcal{F}_{B,1}$ and $u^{*}(\bm{x})$ is the exact solution of Eq. (6), then there is a constant $C$ such that $\mathcal{E}_{D,r}(\phi)-\mathcal{E}_{D,r}(u^{*})\leq C\|\phi-u^{*}\|_{H^{1}(\widehat{\Omega})}$.
@@
The error is decomposed into three components. The term $\left(\frac{W}{(\log W)^{2}}\right)^{-\frac{4(n-1)}{d}}$ represents the optimal approximation error, while $\frac{\log M_{1}}{{M_{1}}}$ corresponds to the sample error in the interior, and $\frac{\log M_{2}}{{M_{2}}}$ denotes the sample error in the boundary.
@@ 4.2 Generalization error of Physics-Informed Neural Network Methods
In this subsection, we apply PINN methods to solve the Poisson equation Eq. (6). PINN differs from the deep Ritz method in that it uses the residual error as the loss function, which is an $H^{2}$ loss function, similar to the DGM. Therefore, we will leverage the result from Theorem 3.2 for the case where $k=2$. The corresponding loss function in Physics-Informed Neural Network methods can be expressed as:
@@
Suppose that $\phi(\bm{x};\bm{\theta}_{D,p})\in\mathcal{F}_{B,2}$ and $u^{*}$ is the exact solution of Eq. (6), then there is a constant $C$ such that $\mathcal{E}_{D,p}(\phi(\bm{x};\bm{\theta}_{D,p}))\leq C\|\phi(\bm{x};\bm{\theta}_{D,p})-u^{*}\|_{H^{2}(\widehat{\Omega})}$.
@@ Appendix B Lemmas Related to Covering Numbers
Denote $v=\phi(\bm{x};\bm{\theta}_{D,r})-u^{*}$, then we have
$\mathcal{E}_{D,r}(\phi(\bm{x};\bm{\theta}_{D,r}))-\mathcal{E}_{D,r}(u^{*})=\mathcal{E}_{D,r}(v+u^{*})-\mathcal{E}_{D,r}(u^{*})$
$=$ $\int_{\widehat{\Omega}}\nabla v\nabla u^{*}\,\mathrm{d}\bm{x}+\frac{1}{2}\int_{\widehat{\Omega}}|\nabla v|^{2}\,\mathrm{d}\bm{x}-\int_{\widehat{\Omega}}fv\,\mathrm{d}\bm{x}+\lambda\int_{\partial\widehat{\Omega}}v^{2}\,\mathrm{d}s.$ (43)
Due to $u^{*}$ being an exact solution and trace theorem inequality, we have
$\int_{\widehat{\Omega}}\nabla v\nabla u^{*}\,\mathrm{d}\bm{x}-\int_{\widehat{\Omega}}fv\,\mathrm{d}\bm{x}=\int_{\partial\widehat{\Omega}}v\frac{\partial u^{*}}{\partial\nu}\,\mathrm{d}\bm{s}\leq C\|u^{*}\|_{H^{2}(\widehat{\Omega})}\|v\|_{H^{1}(\widehat{\Omega})}$.
Therefore,
$\mathcal{E}_{D,r}(\phi(\bm{x};\bm{\theta}_{D,r}))-\mathcal{E}_{D,r}(u^{*})=\frac{1}{2}\int_{\widehat{\Omega}}|\nabla v|^{2}\,\mathrm{d}\bm{x}+\lambda\int_{\partial\widehat{\Omega}}v^{2}\,\mathrm{d}s$ (44)
Finally, due to the trace theorem inequality, we have that
$\frac{1}{2}\int_{\widehat{\Omega}}|\nabla v|^{2}\,\mathrm{d}\bm{x}+\lambda\int_{\partial\widehat{\Omega}}v^{2}\,\mathrm{d}s\leq C\|v(\bm{x})\|_{H^{1}(\widehat{\Omega})}\leq C\|\phi(\bm{x};\bm{\theta}_{D,r})-u^{*}\|_{H^{1}(\widehat{\Omega})}$.
@@
Suppose that $\phi(\bm{x};\bm{\theta}_{D,p})\in\mathcal{F}_{B,2}$ and $u^{*}$ is the exact solution of Eq. (6), then there is a constant $C$ such that
$\mathcal{E}_{D,p}(\phi(\bm{x};\bm{\theta}_{D,p}))\leq C\|\phi(\bm{x};\bm{\theta}_{D,p})-u^{*}\|_{H^{2}(\widehat{\Omega})}$.

### v4

@@ 4.1 Generalization Error in deep Ritz Method
We initially focus on solving the Poisson equation, as represented by Eq. (6) with Neumann boundary condition, using the deep Ritz method. This approach utilizes the energy as the loss function during training to guide the neural networks in approximating the solutions. In the context of the Poisson equation, the energy functional encompasses the first derivative of the neural networks. Consequently, based on (lu2021priori), the corresponding loss function in deep Ritz methods can be expressed as:
$\mathcal{E}_{D,r}(\phi):=$ $\int_{\widehat{\Omega}}g(\bm{x};\bm{\theta})\,\mathrm{d}\bm{x}+\left(\int_{\widehat{\Omega}}\phi(\bm{x};\bm{\theta})\,\mathrm{d}\bm{x}\right)^{2},$
@@
where $\bm{\theta}$ represents all the parameters in the neural network. To discretize the loss function, we randomly choose $M_{1}$ points in the interior and $M_{2}$ points on the boundary:
$\mathcal{E}_{S,r}(\phi):=$ $\frac{1}{2M_{1}}\sum_{i=1}^{M_{1}}g(\bm{x}_{i};\bm{\theta})+\left(\frac{1}{M_{2}}\sum_{i=1}^{M_{2}}\phi(\bm{y}_{i};\bm{\theta})\right)^{2}.$
@@
Suppose that $\phi\in\mathcal{F}_{B,1}$ and $u^{*}(\bm{x})$ is the exact solution of Eq. (6) with Neumann boundary condition, then there is a constant $C$ such that $\mathcal{E}_{D,r}(\phi)-\mathcal{E}_{D,r}(u^{*})\leq C\|\phi-u^{*}\|^{2}_{H^{1}(\widehat{\Omega})}.$
@@
The error is decomposed into three components. The term $\left(\frac{W}{(\log W)^{2}}\right)^{-\frac{4(n-1)}{d}}$ represents the optimal approximation error, while $\frac{\log M_{i}}{{M_{i}}}$ corresponds to the sample error.
@@ 4.2 Generalization error of Physics-Informed Neural Network Methods
In this subsection, we apply PINN methods to solve the Poisson equation Eq. (6) with Dirichlet boundary condition. PINN differs from the deep Ritz method in that it uses the residual error as the loss function, which is an $H^{2}$ loss function, similar to the DGM. Therefore, we will leverage the result from Theorem 3.2 for the case where $k=2$. The corresponding loss function in Physics-Informed Neural Network methods can be expressed as:
@@
Suppose that $\phi(\bm{x};\bm{\theta}_{D,p})\in\mathcal{F}_{B,2}$ and $u^{*}$ is the exact solution of Eq. (6) with Dirichlet boundary condition, then there is a constant $C$ such that $\mathcal{E}_{D,p}(\phi(\bm{x};\bm{\theta}_{D,p}))\leq C\|\phi(\bm{x};\bm{\theta}_{D,p})-u^{*}\|^{2}_{H^{2}(\widehat{\Omega})}.$
@@ Appendix B Lemmas Related to Covering Numbers
The proof can be found in [lu2021priori]. ∎
@@
Suppose that $\phi(\bm{x};\bm{\theta}_{D,p})\in\mathcal{F}_{B,2}$ and $u^{*}$ is the exact solution of Eq. (6) with Dirichlet boundary condition, then there is a constant $C$ such that
$\mathcal{E}_{D,p}(\phi(\bm{x};\bm{\theta}_{D,p}))\leq C\|\phi(\bm{x};\bm{\theta}_{D,p})-u^{*}\|^{2}_{H^{2}(\widehat{\Omega})}.$